# A note on solving basic equations over the semiring of functional digraphs

### v1

@@ Abstract
Endowing the set of functional graphs (FGs) with the sum (disjoint union of graphs) and product (standard direct product on graphs) operations induces on FGs a structure of a commutative semiring $R$. The operations on $R$ can be naturally extended to the set of univariate polynomials $R[X]$ over $R$. This paper provides a polynomial time algorithm for deciding if equations of the type $AX=B$ have solutions when $A$ is just a single cycle and $B$ a set of cycles of identical size. We also prove a similar complexity result for some variants of the previous equation.
@@
functional graphs, direct product, graphs factorization, equations on graphs
@@ 1 Introduction
A functional digraph (FG) is the digraph of a function with finite domain i.e. a digraph with outgoing degree $1$. Endowing the set of functional digraphs FGs with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations provides FGs with the structure of a commutative semiring $R$ in which the empty graph (resp., the single loop) is the neutral element of addition (resp., of product) Dennunzio et al. (2018). This semiring can be naturally extended to the semiring of multivariate polynomials $R[X_{1},X_{2},\ldots]$.
A polynomial in $R[X_{1},X_{2},\ldots]$ represents a (infinite) set of functions which have a common substructure provided by the coefficients (modulo isomorphism). The factorization is an interesting inverse problem in this context. Indeed, assume to have a functional graph $G$. Can $G$ be factozised i.e. decomposed into the product or the sum (or a combination of the two) of functional graphs with smaller vertex set [^1] We stress that factorization here is meant in the context of abstract algebra which has nothing to do with the notion of factor graph often used in graph theory.? If this question has been largely investigated for general graphs, little is known for functional digraphs.
Let us consider another version of the factorization problem. Assume that we have partial information (or partial assumptions) on the factorization of $G$. This partial information is represented by a sequence of functional digraphs $A_{1},A_{2},\ldots,A_{k}$ and the question of factorization can be reformulated as `does the following equation
$A_{1}\cdot X_{1}+A_{2}\cdot X_{2}+\ldots A_{k}\cdot X_{k}=G$
admit a solution?'And more generally one can ask for solutions of the following
$A_{1}\cdot X_{1}^{w_{1}}+A_{2}\cdot X_{2}^{w_{2}}+\ldots A_{k}\cdot X_{k}^{w_{k}}=G$ (1)
@@
where $A$ is made by a single cycle of size $p$ and $B$ is a union of cycles of size $q$ (of course, it might be $p\neq q$). In this paper we prove that deciding if equations of type (2) have solutions or not is in ¶.
The interest of this result is manyfold. Without any doubt it will significantly improve the performances of the software pipeline in  Dennunzio et al. (2023). It will shed some light in some connected important questions concerning the cancellation problem for functional graphs i.e. the problem of establishing if $G\cdot F\approx G\cdot H$ implies $F\approx H$ where $F,G$ and $H$ are FGs and $\approx$ is graph isomorphism Doré et al. (2022); Naquin and Gadouleau (2022). We refer the reader to the classical book of Hammak for the factorization and the cancellation problems for more general graphs Hammack et al. (2011).
@@ 2 Background and basic facts
In Dennunzio et al. (2018), an abstract algebraic setting for studying finite discrete dynamical systems was introduced. This setting, can be adapted in a straightforward way to FGs. In this section we recall just the minimal concepts to understand the problem we want to solve.
@@
As already informally stated in the introduction we have the following algebraic characterisation of the class of FGs.
Theorem 1 (Dennunzio et al. (2018)).
The class of FGs (modulo isomorphisms) equipped with the operations of $+$ and normal-$\cdot$ is a commutative semi-ring.
@@
From Dennunzio et al. (2018) we also recall the following important result about the complexity of solving polynomial equations.
@@ 3 Solving basic equations on permutations
From now on we will focus on a subclass of functional graphs, namely the graphs of permutations. For this reason, we introduce a convenient notation called C-notation taken from Dennunzio et al. (2023). We note $C^{1}_{p}$ the graph made by a single cycle of size $p$. Similarly, $C^{n}_{p}$ denotes a graph which is the sum of $n$ graphs $C^{1}_{p}$. According to this notation, it is clear that $C^{n_{1}}_{p}+C^{n_{2}}_{p}=C^{n_{1}+n_{2}}_{p}$ and $kC^{n}_{p}=C^{kn}_{p}$ for any natural $n_{1},n_{2}$ and $k$. With this notation, the classical result that the graph $G$ of a permutation is a union of disjoint cycles translates into
@@
We proceed by induction over $l$. First of all, we prove that the statement is true for $l=2$, i.e.,
@@
According to the $C$-notation basic equations will have the following form
@@
As we already stressed in the introduction, in many cases, we are not interested in knowing the exact solutions of a basic equation over permutations but we are interested in deciding if an equation admits a solution or not. More formally, in this paper we are interested in the computational complexity of the following decision problem.
@@
$n$ $=n_{1}^{m_{1}}\cdot\ldots\cdot n_{s}^{m_{t}}\text{ with }\operatorname{factors}\left(n\right)=\left\{n_{1},\ldots,n_{t}\right\}\enspace.$
The equation $C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}=C^{n}_{q}$ has solutions if and only if the two following conditions hold
@@
    for all $i\in\left\{1,\ldots,r\right\}$ either $p_{i}^{h_{i}}\in\left\{q_{1}^{k_{1}},\ldots,q_{1}^{k_{s}}\right\}$ or $p_{i}^{h_{i}}$ divides $n$.
@@
Suppose now that for some $i\in\left\{1,\ldots,r\right\}$, ${p_{i}^{h_{i}}}$ is not contained in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, there exists $j\in\left\{1,\ldots,s\right\}$ such that $p_{i}=q_{j}$ and $h_{i}<k_{j}$. Now, if $p_{i}^{h_{i}}$ divides $n$, we have
@@
Then, according to Proposition 1, $p_{i}^{h_{i}}$ (i.e. $q_{j}^{y}$) must divide $n$. ∎
@@
Let $p,q$ be two positive integers. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{s}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{F}$ be the set of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$. Algorithm 1 computes $\Pi_{\mathbf{F}}=\prod_{\mathbf{f}\in\mathbf{F}}\mathbf{f}$ without using the factorizations of $p$ and $q$. Moreover, the time complexity of Algorithm 1 is $O(\log p\cdot\log q)$.
@@
Denote $P=\operatorname{factors}\left(p\right)$ and $Q=\operatorname{factors}\left(q\right)$. If $g=\gcd(p,q)=1$ (line $2$), then $P\cap Q=\emptyset$ and hence $\Pi_{\mathbf{F}}=q$. Hence the algorithm returns $q$ (line $3$). If $g\neq 1$, then taking $q/g$ keeps unchanged the $q_{j}^{k_{j}}$ belonging to $Q\setminus P$ and at the same time decreases of at least $1$ the exponents of the $q_{j}$ belonging to $P\cap Q$. Hence, calling recursively (line $5$) $\Pi_{\mathbf{F}}$ with $(q=q/g,g)$ keeps decreasing the exponents of $q_{j}$ belonging to $P\cap Q$ until they become 00. At that point, $\gcd(q,g)=1$ and the algorithm exits returning $\Pi_{\mathbf{F}}$. Remark that $\gcd(p,q)$ can be computed in $O(\log(\min(p,q)))$ and that the depth of the recursion is $O(\log q)$ in the worst case. Hence, the worst case complexity is $O(\log p\cdot\log q)$. ∎
@@ 2      $g\leftarrow\gcd(p,q)$;
Let $p$ and $q$ be two positive integers such that $p|q$. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{s}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{E}$ be the set of all $p_{i}=q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Then, Algorithm 2 computes ${\Pi}_{\mathbf{E}}=\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$ without using the factorizations of $p$ and $q$. The time complexity of Algorithm 2 is $O(\log p\cdot\log\log p)$.
@@
By executing lines $2$ and $3$ of Algorithm 2 we save in $g$ the product of the factors $q^{k_{j}}_{j}$ such that $q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$. However, the exponents of the factors $q_{j}$ in $g$ are those of $q$ while to compute $\mathbf{\Pi_{E}}$ we need those of $p$. This is the purpose of the while loop (lines $5$-$9$). Indeed, by recursively squaring $d$ it will make the exponents of the factors of $q$ selected in $g$ grow bigger than those in $p$ and hence the $\gcd$ of line $8$ will select the factors and the corresponding exponents that we were looking for. Concerning the complexity, it is enough to remark that the number of iterations of the while loop is $O(\log\log p)$ in the worst case and that each iteration costs $O(\log(\min(q/p,p)))$. ∎
@@ 6                $e\leftarrow\Pi_{\mathbf{E}}(p,\frac{q}{{\Pi}_{\mathbf{F}}(p,q)})$;
Since $p$ divides $q$, we want to calculate ${\Pi}_{\mathbf{E}}(\frac{8316000}{{\Pi}_{\mathbf{F}}(8316000,8400)},8400)$. Let us begin by considering ${\Pi}_{\mathbf{F}}(8316000$ ,8400),8400), 8400 ). Since $8316000$ and $8400$ are not coprime, the method is iterated, i.e.
@@
Once ${\Pi}_{\mathbf{F}}$ has been calculated, through $\frac{q}{{\Pi}_{\mathbf{F}}(q,p)}$ we obtain the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\cap\operatorname{factors}\left(p\right)$, i.e. $2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7$. Then, the goal of ${\Pi}_{\mathbf{E}}$ is to calculate the product of all $q_{j}^{k_{j}}$ such that $q_{j}=p_{i}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Note that $90=2\cdot 3^{2}\cdot 5$.
@@
$C^{n}_{q}=C_{p}^{m}\cdot\sum_{i=1}^{l}C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{p}^{m}\cdot C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{\operatorname{lcm}(p,t_{i})}^{ms_{i}}=C_{q}^{\sum_{i=1}^{l}ms_{i}}=C_{q}^{m\sum_{i=1}^{l}s_{i}}$
and hence Equation (5) has a solution iff $m$ divides $n$ and $C_{p}^{1}\cdot X=C^{n/m}_{q}$ has a solution. We conclude that the time complexity of deciding if Equation (5) has a solution is the max between $O(\log\min(m,n))$ (for checking if $m$ divides $n$) and the time complexity of DEEP with parameters $p,q$ and $n/m$ (see Theorem 4). ∎
@@
Since $\hat{X}$ is a solution it must be either $q=\operatorname{lcm}(p_{k},t_{i})$ or $s_{i}=0$. Hence, from $p_{k}t_{i}=\gcd(p_{k},t_{i})\operatorname{lcm}(p_{k},t_{i})$ and the previous consideration we get $\gcd(p_{k},t_{i})=p_{k}t_{i}/q$. If we replace this last quantity in Equation (8) we get
@@
Let us make some remarks concerning the proof of the previous results. Assume that $H=\frac{nq}{\sum_{k=1}^{r}m_{k}p_{k}}$ is an integer. Then the Diophantine equation
@@
has solutions iff $\gcd(t_{1},\ldots,t_{|d(q)|})=q$ divides $H$ which is always true if $\sum_{k=1}^{r}m_{k}p_{k}$ divides $n$. With standard techniques (see for instance (Niven et al., 1991, Theorem 5.1, page 213)), one can recursively find the solutions in logarithmic time in $q$. We point out that it might happen that all the solutions are non admissible i.e. every solution contains at least one $s_{i}<0$ for $i\in\left\{1,\ldots,r\right\}$. This issue can also be settled in logarithmic time in $q$ (see the considerations following Theorem 5.1 in (Niven et al., 1991, pages 213-214)). However, the big problem that we have is that we need to compute the divisors of $q$ and, as far as we know, the best algorithms for doing that have time complexity which is a polynomial in $q$.
@@
From Equation (8) in the proof of Proposition 2, it is easy to remark that it holds iff
@@
Again, we would like to stress that solutions to Equation (13) can be found in logarithmic time in $\min{m_{1},\ldots,m_{n}}$ giving raise to $r$ equations
@@
()($\Rightarrow$)( ⇒ ) Assume that $\sum_{k=1}^{l}C_{v_{k}}^{u_{k}}$ is a solution for the equation. Then we have
@@
which implies $\operatorname{lcm}(p_{i},v_{k})=q^{t}$. Since $q$ is prime we deduce that either $p_{i}=q^{t}$ or $v_{k}=q^{t}$. Using the hypothesis we have that $v_{k}=q^{t}$ for all $k\in\left\{1,\ldots,l\right\}$. Hence, the solution can be written as $C_{q^{t}}^{u}$. Now, for each $i\in R$ we have $\gcd(p_{i},q^{t})=p_{i}$. Hence, $C_{q^{t}}^{u}$ is a solution iff the following holds
@@
Concerning the uniqueness of the solution, it is enough to remark that question is equivalent to ask for solutions of a Diophantine equation on a single variable. ∎

### v3

# On solving basic equations over the semiring of functional digraphs

@@ Abstract
Endowing the set of functional digraphs (FGs) with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations induces on FGs a structure of a commutative semiring $R$. The operations on $R$ can be naturally extended to the set of univariate polynomials $R[X]$ over $R$. This paper provides a polynomial time algorithm for deciding if equations of the type $AX=B$ have solutions when $A$ is just a single cycle and $B$ a set of cycles of identical size. We also prove a similar complexity result for some variants of the previous equation.
@@
functional digraphs, direct product, graphs factorization, equations on digraphs
@@ 1 Introduction
A functional digraph (FG) is the digraph of a function with finite domain i.e., a digraph with outgoing degree $1$. Endowing the set of functional digraphs FGs with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations provides FGs with the structure of a commutative semiring $R$ in which the empty graph (resp., the single loop) is the neutral element of addition (resp., of product) [3]. This semiring can be naturally extended to the semiring of multivariate polynomials $R[X_{1},X_{2},\ldots,X_{k}]$.
A polynomial in $R[X_{1},X_{2},\ldots,X_{k}]$ represents a (infinite) set of functions which have a common substructure provided by the coefficients (modulo isomorphism). The factorization is an interesting inverse problem in this context. Indeed, assume to have a functional graph $G$. Can $G$ be factorized i.e., decomposed into the product or the sum (or a combination of the two) of functional graphs with smaller vertex set [^1] We stress that factorization here is meant in the context of abstract algebra which has nothing to do with the notion of factor graph often used in graph theory.? If this question has been largely investigated for general graphs, little is known for functional digraphs.
Let us consider another version of the factorization problem. Assume that we have partial information (or partial assumptions) on the factorization of $G$. This partial information is represented by a sequence of functional digraphs $A_{1},A_{2},\ldots,A_{k}$ and the question of factorization can be reformulated as ‘does the following equation
$A_{1}\cdot X_{1}+A_{2}\cdot X_{2}+\ldots+A_{k}\cdot X_{k}=G$
admit a solution?’ And more generally one can ask for solutions of the following
$A_{1}\cdot X_{1}^{w_{1}}+A_{2}\cdot X_{2}^{w_{2}}+\ldots+A_{k}\cdot X_{k}^{w_{k}}=G$ (1)
@@
where $A$ is made by a single cycle of size $p$ and $B$ is a union of cycles of size $q$ (of course, it might be $p\neq q$). In this paper, we prove that deciding if equations of type (2) have solutions or not is in ¶.
The interest of this result is manyfold. Without any doubt it will significantly improve the performances of the software pipeline in  [4] for the case in which one is interested to the existence of solutions and not to their enumeration. It will shed some light in some connected important questions concerning the cancellation problem for functional graphs i.e., the problem of establishing if $G\cdot F\approx G\cdot H$ implies $F\approx H$ where $F,G$ and $H$ are FGs and $\approx$ is graph isomorphism [5, 8]. We refer the reader to the classical book of Hammak for the factorization and the cancellation problems for more general graphs [6].
@@ 2 Background and basic facts
In [3], an abstract algebraic setting for studying finite discrete dynamical systems was introduced. This setting, can be adapted in a straightforward way to FGs. In this section, we recall just the minimal concepts to understand the problem we want to solve.
@@
As already informally stated in the introduction, we have the following algebraic characterisation of the class of FGs.
Theorem 1 ([3]).
The class of FGs (modulo isomorphisms) equipped with the operations of $+$ and $\cdot$ is a commutative semiring.
@@
From [3], we also recall the following important result about the complexity of solving polynomial equations.
@@ 3 Solving basic equations on permutations
From now on, we will focus on a subclass of functional graphs, namely the graphs of permutations. For this reason, we introduce a convenient notation called C-notation taken from [4]. We note $C^{1}_{p}$ the graph made by a single cycle of size $p$. Similarly, $C^{n}_{p}$ denotes a graph which is the sum of $n$ graphs $C^{1}_{p}$. According to this notation, it is clear that $C^{n_{1}}_{p}+C^{n_{2}}_{p}=C^{n_{1}+n_{2}}_{p}$ and $kC^{n}_{p}=C^{kn}_{p}$ for any natural $n_{1},n_{2}$ and $k$. With this notation, the classical result that the graph $G$ of a permutation is a union of disjoint cycles translates into
@@
We proceed by induction over $l$. First of all, we prove that the statement is true for $l=2$, i.e.,,
@@
According to the $C$-notation, basic equations will have the following form
@@
As we already stressed in the introduction, in many cases, we are not interested in knowing the exact solutions of a basic equation over permutations but we are interested in deciding if an equation admits a solution or not. More formally, in this paper, we are interested in the computational complexity of the following decision problem.
@@
$n$ $=n_{1}^{m_{1}}\cdot\ldots\cdot n_{t}^{m_{t}}\text{ with }\operatorname{factors}\left(n\right)=\left\{n_{1},\ldots,n_{t}\right\}\enspace.$
The equation $C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}=C^{n}_{q}$ has solutions if and only if the two following conditions hold:
@@
    for all $i\in\left\{1,\ldots,r\right\}$ either $p_{i}^{h_{i}}\in\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$ or $p_{i}^{h_{i}}$ divides $n$.
@@
Suppose now that, for some $i\in\left\{1,\ldots,r\right\}$, ${p_{i}^{h_{i}}}$ is not contained in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, there exists $j\in\left\{1,\ldots,s\right\}$ such that $p_{i}=q_{j}$ and $h_{i}<k_{j}$. Now, if $p_{i}^{h_{i}}$ divides $n$, we have
@@
Then, according to Proposition 1, $p_{i}^{h_{i}}$ (i.e., $q_{j}^{y}$) must divide $n$. ∎
@@
Let $p,q$ be two positive integers. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{F}$ be the set of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$. Algorithm 1 computes $\Pi_{\mathbf{F}}=\prod_{\mathbf{f}\in\mathbf{F}}\mathbf{f}$ without using the factorizations of $p$ and $q$. Moreover, the time complexity of Algorithm 1 is $O(\log p\cdot\log q)$.
@@
Denote $P=\operatorname{factors}\left(p\right)$ and $Q=\operatorname{factors}\left(q\right)$. If $g=\gcd(p,q)=1$ (Line $2$), then $P\cap Q=\emptyset$ and $\Pi_{\mathbf{F}}=q$.Hence, the algorithm returns $q$ (Line $4$). If $g\neq 1$, then taking $q/g$ keeps unchanged the $q_{j}^{k_{j}}$ belonging to $Q\setminus P$ and at the same time decreases of at least $1$ the exponents of the $q_{j}$ belonging to $P\cap Q$. Therefore, calling recursively (Line $6$) $\Pi_{\mathbf{F}}$ with $(q=q/g,g)$ keeps decreasing the exponents of $q_{j}$ belonging to $P\cap Q$ until they become 00. At that point, $\gcd(q,g)=1$ and the algorithm exits returning $\Pi_{\mathbf{F}}$. Remark that $\gcd(p,q)$ can be computed in $O(\log(\min(p,q)))$ and that the depth of the recursion is $O(\log q)$ in the worst case. Hence, the worst case complexity is $O(\log p\cdot\log q)$. ∎
@@ 2      $g\leftarrow\gcd(p,q)$;
Let $p$ and $q$ be two positive integers such that $p|q$. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{E}$ be the set of all $p_{i}=q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Then, Algorithm 2 computes ${\Pi}_{\mathbf{E}}=\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$ without using the factorizations of $p$ and $q$. The time complexity of Algorithm 2 is $O(\log p\cdot\log\log p)$.
@@
By executing Lines $2$ and $3$ of Algorithm 2, we save in $g$ the product of the factors $q^{k_{j}}_{j}$ such that $q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$. However, the exponents of the factors $q_{j}$ in $g$ are those of $q$ while to compute $\mathbf{\Pi_{E}}$ we need those of $p$. This is the purpose of the while loop (Lines $5$-$9$). Indeed, by recursively squaring $d$, it will make the exponents of the factors of $q$ selected in $g$ grow bigger than those in $p$ and hence the $\gcd$ of line $8$ will select the factors and the corresponding exponents that we were looking for. Concerning the complexity, it is enough to remark that the number of iterations of the while loop is $O(\log\log p)$ in the worst case and that each iteration costs $O(\log(\min(q/p,p)))$. ∎
@@ 6                $e\leftarrow\Pi_{\mathbf{E}}(p,\frac{q}{{\Pi}_{\mathbf{F}}(p,q)})$;
Since $p$ divides $q$, we want to calculate ${\Pi}_{\mathbf{E}}(\frac{8316000}{{\Pi}_{\mathbf{F}}(8316000,8400)},8400)$. Let us begin by considering ${\Pi}_{\mathbf{F}}(8316000$ ,8400),8400), 8400 ). Since $8316000$ and $8400$ are not coprime, the method is iterated, i.e.,
@@
Once ${\Pi}_{\mathbf{F}}$ has been calculated, through $\frac{q}{{\Pi}_{\mathbf{F}}(q,p)}$ we obtain the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\cap\operatorname{factors}\left(p\right)$, i.e., $2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7$. Then, the goal of ${\Pi}_{\mathbf{E}}$ is to calculate the product of all $q_{j}^{k_{j}}$ such that $q_{j}=p_{i}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Note that $90=2\cdot 3^{2}\cdot 5$.
@@
$C^{n}_{q}=C_{p}^{m}\cdot\sum_{i=1}^{l}C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{p}^{m}\cdot C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{\operatorname{lcm}(p,t_{i})}^{ms_{i}\gcd(p,t_{i})}=C_{q}^{\sum_{i=1}^{l}\frac{mq}{p}\cdot\frac{s_{i}}{t_{i}}}=C_{q}^{\frac{mq}{p}\sum_{i=1}^{l}\frac{s_{i}}{t_{i}}}$
and hence Equation (5) has a solution iff $\frac{mq}{p}$ divides $n$ and $C_{p}^{1}\cdot X=C^{\frac{np}{mq}}_{q}$ has a solution. We conclude that the time complexity of deciding if Equation (5) has a solution is the max between $O(\log\min(\frac{mq}{p},n))$ (for checking if $\frac{mq}{p}$ divides $n$) and the time complexity of DEEP with parameters $p,q$ and $\frac{np}{mq}$ (see Theorem 4). ∎
@@
Since $\hat{X}$ is a solution it must be either $q=\operatorname{lcm}(p_{k},t_{i})$ or $s_{i}=0$. Hence, from $p_{k}t_{i}=\gcd(p_{k},t_{i})\operatorname{lcm}(p_{k},t_{i})$ and the previous consideration, we get $\gcd(p_{k},t_{i})=p_{k}t_{i}/q$. If we replace this last quantity in Equation (8) we get
@@
Let us make some remarks concerning the proof of the previous results. Assume that $H=\frac{nq}{\sum_{k=1}^{r}m_{k}p_{k}}$ is an integer. Then, the Diophantine equation
@@
has solutions iff $\gcd(t_{1},\ldots,t_{|d(q)|})=q$ divides $H$, which is always true if $\sum_{k=1}^{r}m_{k}p_{k}$ divides $n$. With standard techniques (see for instance [7, Theorem 5.1, page 213]), one can recursively find the solutions in logarithmic time in $q$. We point out that it might happen that all the solutions are non admissible i.e., every solution contains at least one $s_{i}<0$ for $i\in\left\{1,\ldots,r\right\}$. This issue can also be settled in logarithmic time in $q$ (see the considerations following Theorem 5.1 in [7, pages 213-214]). However, the big problem that we have is that we need to compute the divisors of $q$ and, as far as we know, the best algorithms for doing that have time complexity which is a polynomial in $q$ (see [1] for details).
@@
Denote $d(q)$ the set of divisors of $q$. From Equation (8) in the proof of Proposition 2, it is easy to check that $\hat{X}=\sum_{i=1}^{|d(q)|}C_{t_{i}}^{s_{i}}$ is a solution if and only if
@@
Again, we would like to stress that solutions to Equation (13) can be found in logarithmic time in $\min(m_{1},\ldots,m_{n})$ giving raise to $r$ equations
@@
()($\Rightarrow$)( ⇒ ) Assume that $\sum_{k=1}^{l}C_{v_{k}}^{u_{k}}$ is a solution for the equation. Then, we have
@@
which implies $\operatorname{lcm}(p_{i},v_{k})=q^{t}$. Since $q$ is prime, we deduce that either $p_{i}=q^{t}$ or $v_{k}=q^{t}$. Using the hypothesis we have that $v_{k}=q^{t}$ for all $k\in\left\{1,\ldots,l\right\}$. Hence, the solution can be written as $C_{q^{t}}^{u}$. Now, for each $i\in R$ we have $\gcd(p_{i},q^{t})=p_{i}$. Hence, $C_{q^{t}}^{u}$ is a solution iff the following holds
@@
Concerning the uniqueness of the solution, it is enough to remark that the question is equivalent to ask for solutions of a Diophantine equation on a single variable. ∎